OpenStudy (anonymous):

Find the least common multiple of y^2 – 2y – 15 and y^2 – 7y + 10. (1 point) (y – 2)(y + 3)(y – 5) (y + 2)(y + 3)(y – 5) (y – 2)(y – 3)(y – 5) (y + 2)(y + 3)(y + 5)

OpenStudy (anonymous):

Y^2-2y-15 = y^2-5y+3y-15 and u can make factor of this .. similarly other one : y^2-7y+10 = y^2-5y-2y+10 which gives u two factors in each case : In case I : (Y-5)(y+3) and in Case II = (y-5)(y-2) so LCM of these are : (y-5)(y-2)(y+3)